OpenStudy (anonymous):

(6-2i)^2 simplify to a+bi form

OpenStudy (anonymous):

Do it as you would square a real set of terms.

OpenStudy (anonymous):

12i-4i??

OpenStudy (anonymous):

\[(6-2i)^2=6^2+2*(6)*(-2i)+(-2i)^2\]

OpenStudy (anonymous):

\[36-24i-4=32-24i\]